OpenStudy (anonymous):

The following sequence is geometric. What is the next term? 729, 243, 81, 27 ...

ganeshie8 (ganeshie8):

find the common ratio

ganeshie8 (ganeshie8):

r = 27/81 = 1/3

ganeshie8 (ganeshie8):

to find next term, multiply current term 27 with "r"

ganeshie8 (ganeshie8):

27 * \(\frac{1}{3}\) = ?

OpenStudy (anonymous):

um... i got lost alittle

OpenStudy (asnaseer):

oops - sorry @XHUHX I misread this question - it is NOT the same at all - please forgive me :(

OpenStudy (anonymous):

no problem

OpenStudy (asnaseer):

thx :) - as a penance I can try and help on this one? :)

OpenStudy (anonymous):

alright thanks

OpenStudy (asnaseer):

unless @ganeshie8 wants to continue here - I don't want to step on his/her shoes?

OpenStudy (anonymous):

ok

OpenStudy (asnaseer):

:) ok

OpenStudy (asnaseer):

ok XHUHX - where exactly are you stuck here?

OpenStudy (anonymous):

i couldnt find the commone ratio

OpenStudy (asnaseer):

image the terms of the series were: \(a_1, a_2, a_3, ..., a_n\), then the common ratio is found by dividing any term (except the first term) with it's previous term.

OpenStudy (asnaseer):

so in the series I gave, you could get the common ratio in ANY of the following ways:\[\frac{a_2}{a_1}\text{ OR }\frac{a_3}{a_2}\text{ OR }\frac{a_4}{a_3}\text{ OR }...\frac{a_n}{a_{n-1}}\]

OpenStudy (asnaseer):

all those should give the SAME answer. does that make sense?

OpenStudy (anonymous):

would the numbers go in a2/a1?

OpenStudy (asnaseer):

in your case, the series is: 729, 243, 81, 27, ... so the common ratio can be found from ANY of these ratios:\[\frac{243}{729}\text{ OR }\frac{81}{243}\text{ OR }\frac{27}{81}\]

OpenStudy (anonymous):